OpenStudy (anonymous):

simplify the equation to x^2+y^2=r^2 x^2+y^2+6x-10y+33

OpenStudy (anonymous):

Something is missing!

OpenStudy (anonymous):

all of it equals 0

OpenStudy (anonymous):

You can write in the form \((x-a)^2+(y-b)^2=r^2\) by completing squares. That's \((x^2+6x+9)(y^2-10y+25=1 \implies (x+3)^2+(x-5)^2=1\).

OpenStudy (anonymous):

whew bit it still will not be \[x^2+y^2=r^2\] \[(x+3)^2+(y-5)^2=1\]

myininaya (myininaya):

x^2+6x+9-9+y^2-10y+25-25+33=0 (x+3)^2+(y-5)^2=-33+9+25 (x+3)^2+(y-5)^2=1

OpenStudy (anonymous):

Sorry \((x+3)^2+(y-5)^2=1\) :D

OpenStudy (anonymous):

The vertex is \((-3,5)\). The x-intercept can be found by plugging \(y=0\) in the equation. Doing so gives \((x+3)^2+25=1 \implies (x+3)^2=-24\), but \((x+3)^2 \ge0\). So, There is no x-intercept.